OpenStudy (anonymous):

3(r – 0.3r) – 5= 3(r-0.3r-5) Um, can someone help me solve step, by step

OpenStudy (mathmale):

Why repost the same question, when others have already tried to help you with it? I asked you to perform the operations WITHIN PARENTHESES first. Next, carry out the indicated multiplication by 3. Show all steps of your work if you want feedback and guidance.

OpenStudy (mathmale):

Hint: What is r-0.3r? What is: r -0.3r -------- ????

OpenStudy (anonymous):

3x r-0.3r/-5x3 ? I have no idea what you're asking for :( sorry @mathmale

OpenStudy (mathmale):

The work you must do in this problem is similar to that y ou're doing in another problem right now: Using the distributive property of multiplication. 3(r – 0.3r) – 5= 3(r-0.3r-5) is easier to test for equality (or no equality) if you'd first combine the quantities inside parentheses. That r - 0.3r becomes 1.0r-0.3r=0.7r. Then, on the left side of your equation, you have 3(0.7r) - 5. On the right side you have 3(0.7r-5). Please look at this intermediate equation: 3(0.7r) – 5= 3(0.7r-5) and decide whether or not the equation is true.